Show the solution sets written algebraically and as a union or intersection of intervals. Please show all work so I can see how it is done. Thank you.
4x + 7 < 11 or 1 – x ≤ -2

Answer by rothauserc(4718)

1) 4x + 7 < 11
4x < 4
x < 1
2) 1 - x < or = -2
-x < or = -3
x > or = 3
(-infinity, 1), [3, infinity)